Roots of equation x.x -10cx-11d=0 are a and b and those of x.x -10ax-11b=0 are c and d. Then a+b+c+d=?

Grade:11

1 Answers

Vikas TU
7 years ago
From first eqn.
a+b that is sum of roots => 10c
From second eqn.
c+d = >10a
Therfore,
a+b+c+d =  10(a+c)